OpenStudy (anonymous):

You attend an amusement park with your family. Your parents buy you an all-ride pass for $20, shown as f(x). Instead of getting a pass, your parents decide to pay $4 for each ride they take, shown as g(x). What function shows the correct combination of these two functions to represent the total cost to them of attending the amusement park that day, shown as h(x)?

OpenStudy (anonymous):

help me pleaseee

OpenStudy (anonymous):

these are the choices.... f(x) = 20x, g(x) = 4, h(x) = 20x + 4 f(x) = 20, g(x) = 4, h(x) = 4 + 20 f(x) = 20, g(x) = 4x, h(x) = 4x + 20 f(x) = 20x, g(x) = 4x, h(x) = 20x + 4x